all parts please!4. The zeta function (8) = 2n=ln,s > 1, plays an important role in many areas of math- ematics, especially number theory (it
Answer #1

a) Remember that, using a geometric series, we have \frac{1}{1-xy} = \sum_{n=0}^\infty (xy)^n for |xy|<1 . Therefore,

I = \int_0^1 \int_0^1 \frac{1}{1-xy} \,\text dx\,\text dy
= \int_0^1 \int_0^1 \sum_{n=0}^\infty (xy)^n \,\text dx\,\text dy
= \sum_{n=0}^\infty \int_0^1 \int_0^1 x^n y^n \,\text dx\,\text dy
= \sum_{n=0}^\infty \int_0^1 \left( \frac1{n+1} \right )y^n \,\text dy
= \sum_{n=0}^\infty \left( \frac1{n+1} \right ) \left( \frac1{n+1} \right )
= \sum_{n=0}^\infty \frac{1}{(n+1)^2}
= \sum_{n=1}^\infty \frac{1}{n^2} .

b) Using the substitution, xy = \frac 12(u^2 - v^2) and the integral becomes,

I =\iint_{D} \frac{2}{2-u^2 +v^2} \,\text dv\,\text du

where D is the region

06 0.4 0.2 0.2 0.4 0.6 0.8 1 1.2 -0.2 -0.4 -0.6

c)

I =\iint_{D} \frac{2}{2-u^2+v^2} \,\text dv\,\text du
= \int_0^{\sqrt{2}/2} \int_{-u}^u \frac{2}{2-u^2 +v^2}\,\text dv\,\text du+ \int_{\sqrt{2}/2}^{\sqrt 2} \int_{-u}^u \frac{2}{2-u^2 + v^2}\,\text dv\,\text du
= I_1 + I_2 .